OpenStudy (anonymous):

What is the slope of the line between (–4, 4) and (–1, –2)?

OpenStudy (aripotta):

slope formula:

OpenStudy (aripotta):

do you know what to do?

OpenStudy (anonymous):

no lol

OpenStudy (aripotta):

see where it says y2 and y1?

OpenStudy (aripotta):

that's the y part of the coordinates

OpenStudy (anonymous):

Okay, (-4,4) is a coordinate. Do you know which one is x and which is y?

OpenStudy (anonymous):

Yes -4 is x and 4 is y i think..

OpenStudy (aripotta):

that's right

OpenStudy (aripotta):

so if 4 is y1, then what's y2?

OpenStudy (anonymous):

-1 & -2

OpenStudy (aripotta):

no..

OpenStudy (aripotta):

the second coordinate is (-1,-2), right?

OpenStudy (aripotta):

what's the y part of that coordinate?

OpenStudy (anonymous):

yea wouldnt that be x i and 2

OpenStudy (anonymous):

Okay, if -4 is \[x _{1}\] and 4 is \[y _{1}\] What is \[x _{2}\] And what is \[y_{2}\]

OpenStudy (anonymous):

x2= -1 y2=-2

OpenStudy (aripotta):

that's right, so draw for me what the formula would look like with the numbers put in

OpenStudy (anonymous):

Yep! Okay so look at the attachment that @AriPotta put in. It's the slope formula. Plug in x1, y1, x2, and y2, into the formula.

OpenStudy (anonymous):

-4 x 4/ -1 x -2

OpenStudy (aripotta):

it wouldn't be multiplication, you're subtracting. and it has to be y2 - y1/x2 - x1

OpenStudy (aripotta):

do you see what i'm trying to say? :/

OpenStudy (anonymous):

lol yea i see hold on ill tell u the anwser...

OpenStudy (anonymous):

i got 2.66666 i dont think i was right lol

OpenStudy (aripotta):

how did you get that? can you show me?

OpenStudy (anonymous):

-4-4= -8 -1-2= -3 Divide -8 by -3= 2.666666666

OpenStudy (aripotta):

you have to subtract the y1 from y2. you're subtracting y1 from x1. the two y's are on the top and the two x's are on the bottom

OpenStudy (aripotta):

 see what i mean?